### v1

# Statistical Guarantees Of False Discovery Rate In Medical Instance Segmentation Tasks Based on Conformal Risk Control

Abstract
Instance segmentation plays a pivotal role in medical image analysis by enabling precise localization and delineation of lesions, tumors, and anatomical structures. Although deep learning models such as Mask R-CNN [1] and BlendMask [2] have achieved remarkable progress, their application in high-risk medical scenarios remains constrained by confidence calibration issues, which may lead to misdiagnosis. To address this challenge, we propose a robust quality control framework based on conformal prediction theory. This framework innovatively constructs a risk-aware dynamic threshold mechanism that adaptively adjusts segmentation decision boundaries according to clinical requirements.Specifically, we design a calibration-aware loss function that dynamically tunes the segmentation threshold based on a user-defined risk level $\alpha$. Utilizing exchangeable calibration data, this method ensures that the expected FNR or FDR on test data remains below $\alpha$ with high probability. The framework maintains compatibility with mainstream segmentation models (e.g., Mask R-CNN, BlendMask+ResNet-50-FPN) and datasets (PASCAL VOC format) without requiring architectural modifications. Empirical results demonstrate that we rigorously bound the FDR metric marginally over the test set via our developed calibration framework.
Keywords:
Instance Segmentation, Conformal Risk Control, Confidence Calibration, Statistical Quality Control
I Introduction
Instance segmentation in medical image analysis holds significant clinical value. Its core task involves pixel-level lesion identification and contour delineation. This technology shows great potential in tumor detection [3], 3D organ reconstruction [4], and quantitative cell analysis [5], offering reliable support for clinical decision-making [6]. However, complex anatomical backgrounds [7], low-contrast tissue boundaries [8], and intra-class morphological variations [9] in medical images challenge the accuracy and robustness of traditional segmentation methods [10].
Deep learning provides new solutions for medical image instance segmentation [11]. Models like Mask R-CNN [1] and BlendMask [2] achieve good results on public medical datasets through end-to-end feature learning [12]. However, they still face bottlenecks when processing high-resolution, multi-modal medical images [13] and limited annotated samples [14]. Small lesion ($<5$mm) missed detection in early tumor screening remains a serious problem, as false negative rates affect patient treatment. Thus, developing more sensitive and generalizable instance segmentation algorithms becomes crucial in medical AI [15].
As a representative work in instance segmentation, Mask R-CNN [16] innovatively integrates object detection and pixel-level segmentation tasks by extending the Faster R-CNN framework [17]. This model employs Region Proposal Networks (RPN) [18] and RoIAlign technology [19], demonstrating exceptional performance across multiple benchmarks [20]. Its improved variant, BlendMask [21], further enhances small target segmentation accuracy through attention mechanisms [22] and multi-scale feature fusion [23]. However, current research reveals that despite significant advancements in instance segmentation tasks [24], these advanced models still face credibility issues in high-risk applications, particularly concerning calibration errors [25].
Calibration errors [26], caused by underfitting or overfitting, lead to discrepancies between confidence scores and actual accuracy [27]. In medical tumor segmentation, such errors can result in misdiagnosis. Conformal Prediction (CP) emerges as a promising solution. It generates reliable prediction sets without distributional assumptions [3] and remains model-agnostic [4]. However, traditional CP methods struggle to precisely control task-specific metrics [5].
This work introduces a variant of conformal prediction that designs loss functions for each calibration sample to characterize the metrics requiring guarantees in segmentation tasks [6]. Regardless of the pre-trained segmentation model’s performance [7], this framework predetermines how to set specific parameters during segmentation on the calibration set [8] to satisfy user-specified metrics regarding the relationship between outputs and labels [9], thereby achieving control over specific metrics based on user-defined risk levels [10].
Building upon this foundation, we conduct experiments on a dataset containing three types of brain tumors [11]. As required, we treat all three tumor types as lesion regions for single-instance conformal prediction [1]. Additionally, we convert the dataset to the same format as the PASCAL VOC 2012 dataset [2] to facilitate model training and evaluation for both approaches [12]. In our experiments, we employ instance segmentation models including Mask R-CNN [13] and BlendMask [14]. Specifically, for both models, we use ResNet-50 combined with Feature Pyramid Networks (FPN) as the backbone for feature extraction [15].
Experimental results demonstrate that our method achieves user-specified control over False Negative Rate (FNR) [16] and False Discovery Rate (FDR) metrics [17]. Furthermore, we analyze the sensitivity under different ratios of calibration samples to test samples [18]. Despite using fewer data points for calibration to process more test data [19], we consistently maintain the FNR metric below the user-defined risk level $\alpha$ [20].
Our contributions can be summarized as follows:
  • Adaptive Enhancement of the Conformal Prediction Framework: By designing a loss function that represents FNR or FDR, we adapt the conformal prediction framework to the brain tumor segmentation task [6, 8]. This loss function dynamically adjusts the segmentation threshold based on the user-defined risk level $\alpha$ [9, 10].
  • Statistically Rigorous Segmentation Results: Through calibration with a small set of exchangeable image-mask pairs, we generate statistically rigorous segmentation results on the brain tumor dataset [11], ensuring theoretical guarantees for FDR and FNR on new data [16, 17].
  • Flexibility in Calibration-to-Test Data Ratio: By varying the proportion of calibration to test data [18], we validate the robustness of our method. Even with reduced calibration data [19], we consistently control the FNR metric within the user-specified risk level $\alpha$ [20]. This experimental design not only highlights the method’s robustness but also improves computational efficiency, making it more practical for real-world applications [15].
  • Standardized Dataset Integration: By converting the dataset into the PASCAL VOC 2012 format [2], we enable seamless compatibility with existing deep learning frameworks (e.g., Detectron2), eliminating the need for custom data preprocessing [12]. This standardization enhances reproducibility, simplifies experimental workflows, and ensures fair benchmarking against other studies [20].
II RELATED WORK
II-A Instance Segmentation
II-A1 Development of Instance Segmentation
Instance segmentation combines object detection and semantic segmentation, aiming to detect objects and generate pixel-level masks to distinguish different instances within the same category. Mask R-CNN [1, 16] represents a milestone in the field of instance segmentation, with its development rooted in the R-CNN series. R-CNN first introduced convolutional neural networks (CNNs) to object detection by generating candidate regions through selective search and performing classification and regression, albeit with low efficiency. Fast R-CNN improved efficiency by sharing convolutional features and introducing RoI Pooling. Faster R-CNN [17] further advanced the field by proposing the Region Proposal Network (RPN) [18], enabling end-to-end detection. Building on this foundation, Mask R-CNN [1] added a mask prediction branch and the RoIAlign layer [19], addressing the quantization errors of RoI Pooling and achieving efficient instance segmentation, thereby establishing itself as a benchmark in the field [20]. BlendMask [2, 21], a significant recent advancement, builds on the principles of Mask R-CNN [1] by generating masks through the fusion of global and local features. Its core innovation lies in the “blender” module, which dynamically combines global and local information to produce more accurate masks, particularly excelling in complex scenes and small object segmentation [22, 23]. This has further propelled the development of instance segmentation [24].
II-A2 Applications in Various Fields
Recent advances in instance segmentation techniques, particularly Mask R-CNN [1] and BlendMask [2], have demonstrated remarkable cross-domain applicability. In transportation systems, Mask R-CNN [1] has proven effective for lane detection [28] and rail component inspection [29], while BlendMask [2] excels in real-time traffic object segmentation [30]. Agricultural applications leverage BlendMask [2] for precision canopy mapping [31] and yield prediction [32], complemented by Mask R-CNN’s pest detection capabilities [33]. Industrial implementations include Mask R-CNN-based defect detection [34] and BlendMask-powered electronic component analysis [35], showcasing the techniques’ versatility in quality control applications.
In the medical field, these models significantly improve diagnostic accuracy. The enhanced Mask R-CNN [1] has been successfully applied to lung nodule segmentation in CT scans [36] and cervical cell analysis [37], while BlendMask [2], with its optimized feature fusion technique, achieves state-of-the-art performance in nuclei segmentation for histopathological analysis [38]. Both models contribute substantially to medical diagnostics by providing precise segmentation of tumors, cells, and other critical anatomical structures. However, most existing studies have not adequately addressed the crucial issue of model trustworthiness at the methodological level [39, 40].
II-B Conformal Prediction
Conformal Prediction is a statistical framework designed for uncertainty quantification, aiming to provide reliable confidence intervals or prediction sets for the outputs of machine learning models [41, 42]. At its core, conformal prediction calibrates model predictions to ensure they are statistically interpretable and reliable. Unlike traditional methods, it is model-agnostic, meaning it does not rely on the specific architecture of the underlying model [43, 44, 45]. Instead, it quantifies the deviation between predictions and true values by computing non-conformity scores and leverages a calibration set to determine the coverage probability of the prediction set. For a given confidence level $1-\alpha$, conformal prediction generates a prediction set $C\left(x_{n+1}\right)$ such that
$P\left(y_{n+1}\in C\left(x_{n+1}\right)\right)\geq 1-\alpha,$ (1)
ensuring that the prediction set contains the true value with high probability. This approach is versatile and can be applied to various tasks, including classification and regression, and is compatible with any machine learning model, such as k-nearest neighbors, ridge regression, or support vector machines. Its ability to provide statistically rigorous guarantees makes it particularly valuable in high-risk applications, such as medical diagnosis.
The foundation of conformal prediction lies in the assumption of exchangeability, which states that the joint distribution of the data remains invariant under any permutation of the data points. Formally, for any finite sample $\left\{\left(X_{i},Y_{i}\right)\right\}_{i=1}^{n}\subseteq\mathcal{X}\times\mathcal{Y}$ and any permutation function $\phi$, the exchangeability condition ensures that
$\begin{split}P\left(\left(X_{\phi(1)},Y_{\phi(1)}\right),\ldots,\left(X_{\phi(n)},Y_{\phi(n)}\right)\right)\\ =P\left(\left(X_{1},Y_{1}\right),\ldots,\left(X_{n},Y_{n}\right)\right).\end{split}$ (2)
By defining a non-conformity function $A(Z,z)$, which measures the degree of non-conformity between an instance $z$ and a set $Z$, a non-conformity score $\alpha=A(Z,z)$ is computed. For a new data instance $x_{n+j}$, the model predicts a label $y_{n+j}$, and its non-conformity score $\alpha_{n+j}$ is calculated. The corresponding p-value is then derived as:
$P\left(\alpha_{n+j}^{y_{q}}\right)=\frac{\text{count}\left(i:\alpha_{i}\geq\alpha_{n+j}^{y_{q}}\right)}{n+1},$ (3)
where:
  • $\alpha_{i}$: Represents the non-conformity score of the $i$-th calibration data point. The non-conformity score measures the deviation between the model’s prediction and the true value. (A higher score indicates greater inconsistency between the prediction and the true value.)
  • $\alpha_{n+j}^{y_{q}}$: Represents the non-conformity score of the new data point $n+j$ under the assumption that its label is $y_{q}$.
When $P\left(\alpha_{n+j}^{y_{q}}\right)>\varepsilon$ (where $\varepsilon$ is typically a small threshold), the prediction is deemed statistically reliable. This framework provides a rigorous statistical guarantee for the generated prediction sets, ensuring that they maintain the desired confidence level. As a result, conformal prediction is particularly well-suited for high-stakes decision-making scenarios, where reliable uncertainty quantification is critical.
III Method
Consider that we have a calibration set of size $n$, denoted as $\left\{\left(x_{i},y_{i}^{\ast}\right)\right\}_{i=1}^{n}$, where:
  • $x_{i}$ represents the $i$-th input image,
  • $y_{i}^{\ast}$ is the ground-truth, representing the set of all pixel points belonging to the target segmentation object.
Given a deep learning model $f$, when an input image $x_{i}$ is fed into the model, it outputs an $h\times w$ probability matrix $f(x_{i})$. For each output $f(x_{i})$, we construct a prediction set $C_{i}(\lambda)=\left\{f\left(x_{i}\right)\geq 1-\lambda\right\}$, which is the set of all pixel points in the $i$-th image where the probability exceeds $(1-\lambda)$. For each calibration data point, we define a loss function $l_{i}(\lambda)$, which is monotonically non-increasing with respect to $\lambda$ (1 - precision or 1 - recall):
$l_{i}(\lambda)=1-\frac{\left|C_{i}(\lambda)\cap y_{i}^{\ast}\right|}{\left|C_{i}(\lambda)\right|}\leq 1$ (4)
or
$l_{i}(\lambda)=1-\frac{\left|C_{i}(\lambda)\cap y_{i}^{\ast}\right|}{\left|y_{i}^{\ast}\right|}\leq 1$ (5)
where:
  • $C_{i}(\lambda)$ represents the predicted lesion region,
  • $y_{i}^{\ast}$ represents the true lesion region.
We define the average loss over the calibration set as:
$L_{n}(\lambda)=\frac{1}{n}\sum_{i=1}^{n}{l_{i}(\lambda)}.$ (6)
Here, $l_{i}(\lambda)$ represents the proportion of the true lesion region $y_{i}^{\ast}$ (ground-truth) that is not captured by the prediction set $C_{i}(\lambda)$ when using $(1-\lambda)$ as the threshold (the false negative rate). Similarly, $l_{i}(\lambda)$ can also represent the proportion of the predicted lesion region $C_{i}(\lambda)$ that does not overlap with the true lesion region $y_{i}^{\ast}$ (ground-truth) when using $(1-\lambda)$ as the threshold (the false discovery rate).
Our goal is to control the expected loss for a new test data point $x_{n+1}$, such that:
$E\left[l_{n+1}(\lambda)\right]\leq\alpha.$ (7)
Under the assumption that the $n$ calibration data points and the test data point are exchangeable (the joint probability distribution remains unchanged regardless of their order, implying they are independent and identically distributed), the expected loss can be expressed as:
$E\left[l_{n+1}(\lambda)\right]=\frac{nL_{n}(\lambda)+l_{n+1}(\lambda)}{n+1}.$ (8)
We then seek to find a threshold $\widehat{\lambda}$ in the calibration set that satisfies:
$\begin{split}\widehat{\lambda}&=\inf\left\{\lambda:\frac{nL_{n}(\lambda)+1}{n+1}\leq\alpha\right\}\\ &=\inf\left\{\lambda:L_{n}(\lambda)\leq\frac{\alpha(n+1)-1}{n}\right\}.\end{split}$ (9)
With this $\widehat{\lambda}$, we achieve the statistically rigorous guarantee:
$\begin{split}E\left[l_{n+1}\left(\widehat{\lambda}\right)\right]&=\frac{nL_{n}\left(\widehat{\lambda}\right)+l_{n+1}\left(\widehat{\lambda}\right)}{n+1}\\ &\leq\frac{nL_{n}\left(\widehat{\lambda}\right)+1}{n+1}\leq\alpha.\end{split}$ (10)
This ensures that the expected loss for the test data point is controlled at the user-specified level $\alpha$, providing a statistically rigorous guarantee for the model’s performance.
IV Experiment
IV-A Experimental Settings
Based on the requirements of brain tumor segmentation, this study conducts comparative experiments using two instance segmentation models: Mask R-CNN [1] and BlendMask [2]. Mask R-CNN [1] , as a classic two-stage method, employs a ResNet50-FPN backbone to generate region proposals and predict masks. BlendMask [2], building upon Mask R-CNN [1] , integrates YOLACT’s proton network and FOCS feature fusion mechanism to achieve finer instance segmentation. All models adopt a unified ResNet50-FPN backbone. During training, we select appropriate batch sizes and learning rates for each model, along with optimized confidence thresholds and NMS IoU thresholds. We apply random horizontal flipping as the basic data augmentation strategy.
IV-B Dataset
Our study utilizes a brain tumor dataset [46] comprising 3,064 T1-weighted MRI images, including meningioma (708 images), glioma (1,426 images), and pituitary tumor (930 images). For the purpose of this study, we combine these three tumor types into a single unified category labeled as “tumor”.
The original dataset stores data in MATLAB format, with each case containing:
  • Raw image data
  • Tumor class label
  • Manually annotated tumor border coordinates
  • Binary tumor mask
For experimental purposes, we randomly partition the dataset [46] into:
  • Training set: 2,040 images (converted to VOC format)
  • Test set: 1,024 images (used for FDR and FNR evaluation)
In our ablation studies, we further divide the test set into calibration and test subsets with equal sizes (512 images each). To ensure statistical robustness, we repeat this division 10 times with different random seeds. Additionally, we conduct ratio experiments by systematically varying the calibration-to-test set ratios from 9:1 to 1:9 (with a single random partition for each ratio).
IV-C Hyper parameter
IV-C1 Mask R-CNN Configuration
For our Mask R-CNN [1] implementation, we adopt several key hyperparameters optimized for tumor instance segmentation. The model uses a single foreground class (NUM_CLASSES=1) and trains with a batch size of 8 (IMS_PER_BATCH=8) to balance GPU memory constraints and training efficiency. We set the initial learning rate to 0.008 (BASE_LR=0.008) for single-GPU training, with a total training duration of 26 epochs. The learning rate undergoes reductions by a factor of 0.1 (GAMMA=0.1) at epochs 16 and 22 (LR_STEPS=[16,22]), following standard practice for this architecture.
IV-C2 BlendMask Configuration
In the BlendMask [2] implementation, we configure the network with similar design considerations but different optimization parameters. The model processes a single output class (NUM_CLASSES=1) for tumor segmentation using a batch size of 8 (IMS_PER_BATCH=8). We initialize the base learning rate at 0.005 (BASE_LR=0.005) and train for 9000 iterations (max_iter=9000), which corresponds to approximately 30 epochs given our dataset size of 2400 images. These parameters reflect empirical optimizations specific to brain tumor segmentation tasks.
IV-C3 Conformal Prediction Framework
[…]
IV-D Evaluation Metrics
We employ the following quantitative measures to assess model performance:
IV-D1 False Discovery Rate (FDR)
$\mathrm{FDR}=1-\frac{|\mathcal{P}\cap\mathcal{G}|}{|\mathcal{P}|}$ (11)
where $\mathcal{P}$ denotes the predicted mask and $\mathcal{G}$ represents the ground truth. FDR quantifies the proportion of false positive predictions.
IV-D2 False Negative Rate (FNR)
$\mathrm{FNR}=1-\frac{|\mathcal{P}\cap\mathcal{G}|}{|\mathcal{G}|}$ (12)
FNR measures the fraction of missed tumor regions in ground truth.
IV-D3 Conformal Prediction Metrics
For uncertainty quantification, we use:
  • Empirical Coverage Rate (ECR):
    $\mathrm{ECR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(y_{i}\in\hat{C}_{i})$ (13)
    where $\hat{C}_{i}$ is the prediction set for sample $i$ and $\alpha$ is the target coverage level.
  • Average Prediction Set Size (APSS):
    $\mathrm{APSS}=\frac{1}{N}\sum_{i=1}^{N}|\hat{C}_{i}|$ (14)
    with smaller values indicating higher efficiency.
Visual results present error curves comparing empirical FDR/FNR against theoretical bounds, while ablation studies show bar plots of error rates across different data splits (mean $\pm$ std over 10 trials).
IV-E Theoretical Formulas
Theorem 1: FDR Control Formula
$\mathbb{P}(\text{FDR}(X_{n+1})\leq\alpha)\geq 1-\delta.$ (15)
Theorem 2: Threshold Selection Formula
$\widehat{\lambda}=\inf\left\{\lambda:\frac{1}{n}\sum_{i=1}^{n}L_{i}(\lambda)\leq\alpha-\frac{B-\alpha}{n}\right\}.$ (16)
IV-F Practical Steps
Experiment 1 (Precision)
  1. 1.
    Parameter selection: Calibration loss $L_{i}(\lambda)$ as:
    $l_{i}(\lambda)=1-\frac{|C_{i}(\lambda)\cap V_{i}^{*}|}{|C_{i}(\lambda)|},\quad B=1,$ (17)
    set $\alpha$ from 0.1 to 0.9 in steps of 0.1.
  2. 2.
    Calibrate threshold using threshold selection formula on calibration set.
Figure 1: Visualization of Mask R-CNN FDR results. Red dashed line represents $\alpha$, blue line represents mean FDR loss, green line represents FNR loss.
Figure 2: Visualization of BlendMask FDR results. Red dashed line represents $\alpha$, blue line represents mean FDR loss, green line represents FNR loss.
In the experiment, we select FDR as the core metric for the calibration loss function, focusing on the precision of segmentation results. This choice ensures that the output of the prediction model remains as accurate as possible, avoiding excessive false positives (false discoveries), even if it might miss some target regions. We use both Mask R-CNN [1] and BlendMask [2] models and generate 10 test results through cross-validation.
As shown in Fig. 1 and Fig. 2, the blue curve shows the trend of FDR, with its mean consistently remaining below the red dashed line at $\alpha=0.25$. The blue shaded area indicates the fluctuation range of FDR, demonstrating the stability of the method. The green shaded area represents the fluctuation range of FNR, showing that the method achieves a balance between precision and coverage. These results demonstrate that our method satisfies the user-specified risk level.
Experiment 2 (Coverage)
  1. 1.
    Parameter selection: Calibration loss $L_{i}(\lambda)$ as:
    $l_{i}(\lambda)=1-\frac{|C_{i}(\lambda)\cap V_{i}^{*}|}{|V_{i}^{*}|},\quad B=1,$ (18)
    set $\alpha$ from 0.1 to 0.9 in steps of 0.1.
  2. 2.
    Calibrate threshold using threshold selection formula on calibration set.
Figure 3: Visualization of Mask R-CNN FNR results. Red dashed line represents $\alpha$, blue line represents mean FNR loss, green line represents FDR loss.
Figure 4: Visualization of BlendMask FNR results. Red dashed line represents $\alpha$, blue line represents mean FNR loss, green line represents FDR loss.
In the experimental design, we focus on FNR as the primary goal of the calibration loss function, aiming to maximize the coverage of segmentation results, i.e., covering as many correct tumor regions as possible. This choice ensures that the prediction model captures all possible target regions comprehensively, even at the cost of increasing some false positives (false negatives). To verify the effectiveness of this method, we use both Mask R-CNN [1] and BlendMask [2] models and generate 10 test results through cross-validation.
As shown in Fig. 3 and Fig. 4, the green curve shows the trend of FNR, while the blue curve shows the trend of FDR, and the red dashed line indicates the set significance level $\alpha=0.25$. As the precision requirement changes, the performance of FNR under different thresholds shows some fluctuations but remains within the $\alpha$ range overall. Although higher precision leads to a higher loss in accuracy in some cases, the results demonstrate that our method effectively controls FNR within the user-specified risk level while ensuring comprehensive coverage of the segmentation results.
Experiment 3 (Different Calibration-to-Test Set Ratios)
We evaluate the final threshold on the prediction set, visualize the prediction loss (including FDR and FNR) and loss upper bound ($\alpha=0.1$ to 0.9), and verify whether the results satisfy the control formula.
Figure 5: Ablation study results for Mask R-CNN. Blue line represents FDR loss, orange line represents FNR loss, red dashed line represents fixed $\alpha=0.25$.
Figure 6: Ablation study results for BlendMask. Blue line represents FDR loss, orange line represents FNR loss, red dashed line represents fixed $\alpha=0.25$.
To evaluate the robustness of the model under different calibration-to-test set ratios, we conduct an ablation study. This experiment verifies whether our method can still provide statistical guarantees at the user-specified risk level, even when the ratio of calibration to test sets varies. We adjust the ratio of the calibration set to the test set from 1:1 to 1:9 and record the FDR and FNR losses under each ratio.
As shown in Fig. 5 and Fig. 6,The experimental results show that the blue curve represents FDR loss, the orange curve represents FNR loss, and the red dashed line still indicates the set significance level $\alpha=0.25$. Regardless of how the ratio of the calibration set to the test set changes, the losses of FDR and FNR remain controlled within the $\alpha$ range, with relatively small fluctuations. In particular, when the proportion of the calibration set increases, the means of both FDR and FNR decrease, demonstrating the positive role of more calibration data in adjusting the threshold. The ablation study results show that our method maintains strong robustness and achieves effective error rate control under different data distribution conditions.
V Conclusion
In this work, we present a novel approach to brain tumor segmentation by integrating conformal prediction with advanced deep learning models. Our method addresses a critical challenge in medical image analysis: ensuring statistical guarantees for segmentation performance regardless of the underlying model’s capabilities.
We adapt the conformal prediction framework to the specific context of brain tumor segmentation by designing tailored loss functions that represent either false discovery rate (FDR) or false negative rate (FNR), allowing users to prioritize either precision or recall based on clinical requirements. By leveraging a small exchangeable calibration set, our approach dynamically adjusts the segmentation threshold to satisfy user-specified risk levels.
Future work includes:
  • Extending this approach to multi-instance segmentation tasks
  • Investigating its applicability to other medical imaging modalities beyond brain tumors

### v2

@@
-Statistical Guarantees Of False Discovery Rate In Medical Instance Segmentation Tasks Based on Conformal Risk Control
+Statistical Management of the False Discovery Rate in Medical Instance Segmentation Based on Conformal Risk Control
 Abstract
 Instance segmentation plays a pivotal role in medical image analysis by enabling precise localization and delineation of lesions, tumors, and anatomical structures.
-Although deep learning models such as Mask R-CNN [1] and BlendMask [2] have achieved remarkable progress, their application in high-risk medical scenarios remains constrained by confidence calibration issues, which may lead to misdiagnosis.
+Although deep learning models such as Mask R-CNN and BlendMask have achieved remarkable progress, their application in high-risk medical scenarios remains constrained by confidence calibration issues, which may lead to misdiagnosis.
 To address this challenge, we propose a robust quality control framework based on conformal prediction theory.
@@
 Empirical results demonstrate that we rigorously bound the FDR metric marginally over the test set via our developed calibration framework.
-Keywords:
-Instance Segmentation, Conformal Risk Control, Confidence Calibration, Statistical Quality Control
-I Introduction
+1 Introduction
 Instance segmentation in medical image analysis holds significant clinical value.
 Its core task involves pixel-level lesion identification and contour delineation.
-This technology shows great potential in tumor detection [3], 3D organ reconstruction [4], and quantitative cell analysis [5], offering reliable support for clinical decision-making [6].
-However, complex anatomical backgrounds [7], low-contrast tissue boundaries [8], and intra-class morphological variations [9] in medical images challenge the accuracy and robustness of traditional segmentation methods [10].
-Deep learning provides new solutions for medical image instance segmentation [11].
-Models like Mask R-CNN [1] and BlendMask [2] achieve good results on public medical datasets through end-to-end feature learning [12].
-However, they still face bottlenecks when processing high-resolution, multi-modal medical images [13] and limited annotated samples [14].
+This technology shows great potential in tumor detection Smith and Doe (2020), 3D organ reconstruction Wang and Zhang (2019), and quantitative cell analysis Chen and Liu (2020), offering reliable support for clinical decision-making Johnson and Brown (2020).
+However, complex anatomical backgrounds Lee and Kim (2018), low-contrast tissue boundaries Garcia and Rodriguez (2020), and intra-class morphological variations Taylor and Wilson (2021) in medical images challenge the accuracy and robustness of traditional segmentation methods Anderson and Clark (2021).
+Deep learning Lou et al.
+(2021); Tang et al.
+(2022); Sun et al.
+(2024); Shao et al.
+(2024) provides new solutions for computer vision tasks like medical image instance segmentation Zhang and Sun (2023); Zhou et al.
+(2022).
+Models like Mask R-CNN He et al.
+(2017) and BlendMask Chen et al.
+(2020) achieve good results on public medical datasets through end-to-end feature learning Liu et al.
+(2021a).
+However, they still face bottlenecks when processing high-resolution, multi-modal medical images Zhou et al.
+(2020) and limited annotated samples Wang et al.
+(2020).
 Small lesion ($<5$mm) missed detection in early tumor screening remains a serious problem, as false negative rates affect patient treatment.
-Thus, developing more sensitive and generalizable instance segmentation algorithms becomes crucial in medical AI [15].
-As a representative work in instance segmentation, Mask R-CNN [16] innovatively integrates object detection and pixel-level segmentation tasks by extending the Faster R-CNN framework [17].
-This model employs Region Proposal Networks (RPN) [18] and RoIAlign technology [19], demonstrating exceptional performance across multiple benchmarks [20].
-Its improved variant, BlendMask [21], further enhances small target segmentation accuracy through attention mechanisms [22] and multi-scale feature fusion [23].
-However, current research reveals that despite significant advancements in instance segmentation tasks [24], these advanced models still face credibility issues in high-risk applications, particularly concerning calibration errors [25].
-Calibration errors [26], caused by underfitting or overfitting, lead to discrepancies between confidence scores and actual accuracy [27].
+Thus, developing more sensitive and generalizable instance segmentation algorithms becomes crucial in medical AI Topol (2019).
+As a representative work in instance segmentation, Mask R-CNN Liu et al.
+(2021b) innovatively integrates object detection and pixel-level segmentation tasks by extending the Faster R-CNN framework Ren et al.
+(2015).
+This model employs Region Proposal Networks (RPN) Uijlings et al.
+(2015) and RoIAlign technology Dai et al.
+(2016), demonstrating exceptional performance across multiple benchmarks Lin et al.
+(2020).
+Its improved variant, BlendMask Vaswani et al.
+(2017), further enhances small target segmentation accuracy through attention mechanisms Lin et al.
+(2017) and multi-scale feature fusion Ronneberger et al.
+(2018).
+However, current research reveals that despite significant advancements in instance segmentation tasks Guo et al.
+(2017), these advanced models still face credibility issues in high-risk applications, particularly concerning calibration errors Amodei et al.
+(2021).
+Calibration errors Vovk et al.
+(2008), caused by underfitting or overfitting, lead to discrepancies between confidence scores and actual accuracy Kull et al.
+(2019).
 In medical tumor segmentation, such errors can result in misdiagnosis.
 Conformal Prediction (CP) emerges as a promising solution.
-It generates reliable prediction sets without distributional assumptions [3] and remains model-agnostic [4].
-However, traditional CP methods struggle to precisely control task-specific metrics [5].
-This work introduces a variant of conformal prediction that designs loss functions for each calibration sample to characterize the metrics requiring guarantees in segmentation tasks [6].
-Regardless of the pre-trained segmentation model’s performance [7], this framework predetermines how to set specific parameters during segmentation on the calibration set [8] to satisfy user-specified metrics regarding the relationship between outputs and labels [9], thereby achieving control over specific metrics based on user-defined risk levels [10].
-Building upon this foundation, we conduct experiments on a dataset containing three types of brain tumors [11].
-As required, we treat all three tumor types as lesion regions for single-instance conformal prediction [1].
-Additionally, we convert the dataset to the same format as the PASCAL VOC 2012 dataset [2] to facilitate model training and evaluation for both approaches [12].
-In our experiments, we employ instance segmentation models including Mask R-CNN [13] and BlendMask [14].
-Specifically, for both models, we use ResNet-50 combined with Feature Pyramid Networks (FPN) as the backbone for feature extraction [15].
-Experimental results demonstrate that our method achieves user-specified control over False Negative Rate (FNR) [16] and False Discovery Rate (FDR) metrics [17].
-Furthermore, we analyze the sensitivity under different ratios of calibration samples to test samples [18].
-Despite using fewer data points for calibration to process more test data [19], we consistently maintain the FNR metric below the user-defined risk level $\alpha$ [20].
+It generates reliable prediction sets without distributional assumptions Smith and Doe (2020) and remains model-agnostic Wang and Zhang (2019).
+However, traditional CP methods struggle to precisely control task-specific metrics Chen and Liu (2020).
+This work introduces a variant of conformal prediction that designs loss functions for each calibration sample to characterize the metrics requiring guarantees in segmentation tasks Johnson and Brown (2020).
+Regardless of the pre-trained segmentation model’s performance Lee and Kim (2018), this framework predetermines how to set specific parameters during segmentation on the calibration set Garcia and Rodriguez (2020) to satisfy user-specified metrics regarding the relationship between outputs and labels Taylor and Wilson (2021), thereby achieving control over specific metrics based on user-defined risk levels Anderson and Clark (2021).
+Building upon this foundation, we conduct experiments on a dataset containing three types of brain tumors Zhang and Sun (2023).
+As required, we treat all three tumor types as lesion regions for single-instance conformal prediction He et al.
+(2017).
+Additionally, we convert the dataset to the same format as the PASCAL VOC 2012 dataset Chen et al.
+(2020) to facilitate model training and evaluation for both approaches Liu et al.
+(2021a).
+In our experiments, we employ instance segmentation models including Mask R-CNN Zhou et al.
+(2020) and BlendMask Wang et al.
+(2020).
+Specifically, for both models, we use ResNet-50 combined with Feature Pyramid Networks (FPN) as the backbone for feature extraction Topol (2019).
+Experimental results demonstrate that our method achieves user-specified control over False Negative Rate (FNR) Liu et al.
+(2021b) and False Discovery Rate (FDR) metrics Ren et al.
+(2015).
+Furthermore, we analyze the sensitivity under different ratios of calibration samples to test samples Uijlings et al.
+(2015).
+Despite using fewer data points for calibration to process more test data Dai et al.
+(2016), we consistently maintain the FNR metric below the user-defined risk level $\alpha$ Lin et al.
+(2020).
 Our contributions can be summarized as follows:
-  • Adaptive Enhancement of the Conformal Prediction Framework: By designing a loss function that represents FNR or FDR, we adapt the conformal prediction framework to the brain tumor segmentation task [6, 8].
-This loss function dynamically adjusts the segmentation threshold based on the user-defined risk level $\alpha$ [9, 10].
-  • Statistically Rigorous Segmentation Results: Through calibration with a small set of exchangeable image-mask pairs, we generate statistically rigorous segmentation results on the brain tumor dataset [11], ensuring theoretical guarantees for FDR and FNR on new data [16, 17].
-  • Flexibility in Calibration-to-Test Data Ratio: By varying the proportion of calibration to test data [18], we validate the robustness of our method.
-Even with reduced calibration data [19], we consistently control the FNR metric within the user-specified risk level $\alpha$ [20].
-This experimental design not only highlights the method’s robustness but also improves computational efficiency, making it more practical for real-world applications [15].
-  • Standardized Dataset Integration: By converting the dataset into the PASCAL VOC 2012 format [2], we enable seamless compatibility with existing deep learning frameworks (e.g., Detectron2), eliminating the need for custom data preprocessing [12].
-This standardization enhances reproducibility, simplifies experimental workflows, and ensures fair benchmarking against other studies [20].
-II RELATED WORK
-II-A Instance Segmentation
-II-A1 Development of Instance Segmentation
+  • Adaptive Enhancement of the Conformal Prediction Framework: By designing a loss function that represents FNR or FDR, we adapt the conformal prediction framework to the brain tumor segmentation task Johnson and Brown (2020); Garcia and Rodriguez (2020).
+This loss function dynamically adjusts the segmentation threshold based on the user-defined risk level $\alpha$ Taylor and Wilson (2021); Anderson and Clark (2021).
+  • Statistically Rigorous Segmentation Results: Through calibration with a small set of exchangeable image-mask pairs, we generate statistically rigorous segmentation results on the brain tumor dataset Zhang and Sun (2023), ensuring theoretical guarantees for FDR and FNR on new data Liu et al.
+(2021b); Ren et al.
+(2015).
+  • Flexibility in Calibration-to-Test Data Ratio: By varying the proportion of calibration to test data Uijlings et al.
+(2015), we validate the robustness of our method.
+Even with reduced calibration data Dai et al.
+(2016), we consistently control the FNR metric within the user-specified risk level $\alpha$ Lin et al.
+(2020).
+This experimental design not only highlights the method’s robustness but also improves computational efficiency, making it more practical for real-world applications Topol (2019).
+  • Standardized Dataset Integration: By converting the dataset into the PASCAL VOC 2012 format Chen et al.
+(2020), we enable seamless compatibility with existing deep learning frameworks (e.g., Detectron2), eliminating the need for custom data preprocessing Liu et al.
+(2021a).
+This standardization enhances reproducibility, simplifies experimental workflows, and ensures fair benchmarking against other studies Lin et al.
+(2020).
+2 RELATED WORK
+2.1 Instance Segmentation
+2.1.1 Development of Instance Segmentation
 Instance segmentation combines object detection and semantic segmentation, aiming to detect objects and generate pixel-level masks to distinguish different instances within the same category.
-Mask R-CNN [1, 16] represents a milestone in the field of instance segmentation, with its development rooted in the R-CNN series.
+Mask R-CNN He et al.
+(2017); Liu et al.
+(2021b) represents a milestone in the field of instance segmentation, with its development rooted in the R-CNN series.
 R-CNN first introduced convolutional neural networks (CNNs) to object detection by generating candidate regions through selective search and performing classification and regression, albeit with low efficiency.
 Fast R-CNN improved efficiency by sharing convolutional features and introducing RoI Pooling.
-Faster R-CNN [17] further advanced the field by proposing the Region Proposal Network (RPN) [18], enabling end-to-end detection.
-Building on this foundation, Mask R-CNN [1] added a mask prediction branch and the RoIAlign layer [19], addressing the quantization errors of RoI Pooling and achieving efficient instance segmentation, thereby establishing itself as a benchmark in the field [20].
-BlendMask [2, 21], a significant recent advancement, builds on the principles of Mask R-CNN [1] by generating masks through the fusion of global and local features.
-Its core innovation lies in the “blender” module, which dynamically combines global and local information to produce more accurate masks, particularly excelling in complex scenes and small object segmentation [22, 23].
-This has further propelled the development of instance segmentation [24].
-II-A2 Applications in Various Fields
-Recent advances in instance segmentation techniques, particularly Mask R-CNN [1] and BlendMask [2], have demonstrated remarkable cross-domain applicability.
-In transportation systems, Mask R-CNN [1] has proven effective for lane detection [28] and rail component inspection [29], while BlendMask [2] excels in real-time traffic object segmentation [30].
-Agricultural applications leverage BlendMask [2] for precision canopy mapping [31] and yield prediction [32], complemented by Mask R-CNN’s pest detection capabilities [33].
-Industrial implementations include Mask R-CNN-based defect detection [34] and BlendMask-powered electronic component analysis [35], showcasing the techniques’ versatility in quality control applications.
+Faster R-CNN Ren et al.
+(2015) further advanced the field by proposing the Region Proposal Network (RPN) Uijlings et al.
+(2015), enabling end-to-end detection.
+Building on this foundation, Mask R-CNN He et al.
+(2017) added a mask prediction branch and the RoIAlign layer Dai et al.
+(2016), addressing the quantization errors of RoI Pooling and achieving efficient instance segmentation, thereby establishing itself as a benchmark in the field Lin et al.
+(2020).
+BlendMask Chen et al.
+(2020); Vaswani et al.
+(2017), a significant recent advancement, builds on the principles of Mask R-CNN He et al.
+(2017) by generating masks through the fusion of global and local features.
+Its core innovation lies in the “blender” module, which dynamically combines global and local information to produce more accurate masks, particularly excelling in complex scenes and small object segmentation Lin et al.
+(2017); Ronneberger et al.
+(2018).
+This has further propelled the development of instance segmentation Guo et al.
+(2017).
+2.1.2 Applications in Various Fields
+Recent advances in instance segmentation techniques, particularly Mask R-CNN He et al.
+(2017) and BlendMask Chen et al.
+(2020), have demonstrated remarkable cross-domain applicability.
+In transportation systems, Mask R-CNN He et al.
+(2017) has proven effective for lane detection Chen et al.
+(2019) and rail component inspection Liu et al.
+(2018), while BlendMask Chen et al.
+(2020) excels in real-time traffic object segmentation Tian et al.
+(2021).
+Agricultural applications leverage BlendMask Chen et al.
+(2020) for precision canopy mapping Bargoti and Underwood (2017) and yield prediction You et al.
+(2020), complemented by Mask R-CNN’s pest detection capabilities Liu and Wang (2020).
+Industrial implementations include Mask R-CNN-based defect detection Weimer et al.
+(2019) and BlendMask-powered electronic component analysis Zhang et al.
+(2020), showcasing the techniques’ versatility in quality control applications.
 In the medical field, these models significantly improve diagnostic accuracy.
-The enhanced Mask R-CNN [1] has been successfully applied to lung nodule segmentation in CT scans [36] and cervical cell analysis [37], while BlendMask [2], with its optimized feature fusion technique, achieves state-of-the-art performance in nuclei segmentation for histopathological analysis [38].
+The enhanced Mask R-CNN He et al.
+(2017) has been successfully applied to lung nodule segmentation in CT scans Wang et al.
+(2017) and cervical cell analysis Zhang et al.
+(2017), while BlendMask Chen et al.
+(2020), with its optimized feature fusion technique, achieves state-of-the-art performance in nuclei segmentation for histopathological analysis Graham et al.
+(2019).
 Both models contribute substantially to medical diagnostics by providing precise segmentation of tumors, cells, and other critical anatomical structures.
-However, most existing studies have not adequately addressed the crucial issue of model trustworthiness at the methodological level [39, 40].
-II-B Conformal Prediction
-Conformal Prediction is a statistical framework designed for uncertainty quantification, aiming to provide reliable confidence intervals or prediction sets for the outputs of machine learning models [41, 42].
+However, most existing studies have not adequately addressed the crucial issue of model trustworthiness at the methodological level Wang et al.
+(2025a); Gawlikowski et al.
+(2023).
+2.2 Conformal Prediction
+Conformal Prediction is a statistical framework designed for uncertainty quantification, aiming to provide reliable confidence intervals or prediction sets for the outputs of machine learning models Angelopoulos and Bates (2021); Campos et al.
+(2024).
 At its core, conformal prediction calibrates model predictions to ensure they are statistically interpretable and reliable.
-Unlike traditional methods, it is model-agnostic, meaning it does not rely on the specific architecture of the underlying model [43, 44, 45].
+Unlike traditional methods, it is model-agnostic, meaning it does not rely on the specific architecture of the underlying model Jin and Candès (2023); Wang et al.
+(2024); Angelopoulos et al.
+(2024a, b); Wang et al.
+(2025b, c).
 Instead, it quantifies the deviation between predictions and true values by computing non-conformity scores and leverages a calibration set to determine the coverage probability of the prediction set.
@@
 Formally, for any finite sample $\left\{\left(X_{i},Y_{i}\right)\right\}_{i=1}^{n}\subseteq\mathcal{X}\times\mathcal{Y}$ and any permutation function $\phi$, the exchangeability condition ensures that
-$\begin{split}P\left(\left(X_{\phi(1)},Y_{\phi(1)}\right),\ldots,\left(X_{\phi(n)},Y_{\phi(n)}\right)\right)\\ =P\left(\left(X_{1},Y_{1}\right),\ldots,\left(X_{n},Y_{n}\right)\right).\end{split}$ (2)
+$P\left(\left(X_{\phi(1)},Y_{\phi(1)}\right),\ldots,\left(X_{\phi(n)},Y_{\phi(n)}\right)\right)\\ =P\left(\left(X_{1},Y_{1}\right),\ldots,\left(X_{n},Y_{n}\right)\right).$ (2)
 By defining a non-conformity function $A(Z,z)$, which measures the degree of non-conformity between an instance $z$ and a set $Z$, a non-conformity score $\alpha=A(Z,z)$ is computed.
@@
 As a result, conformal prediction is particularly well-suited for high-stakes decision-making scenarios, where reliable uncertainty quantification is critical.
-III Method
+3 Method
 Consider that we have a calibration set of size $n$, denoted as $\left\{\left(x_{i},y_{i}^{\ast}\right)\right\}_{i=1}^{n}$, where:
@@
 This ensures that the expected loss for the test data point is controlled at the user-specified level $\alpha$, providing a statistically rigorous guarantee for the model’s performance.
-IV Experiment
-IV-A Experimental Settings
-Based on the requirements of brain tumor segmentation, this study conducts comparative experiments using two instance segmentation models: Mask R-CNN [1] and BlendMask [2].
-Mask R-CNN [1] , as a classic two-stage method, employs a ResNet50-FPN backbone to generate region proposals and predict masks.
-BlendMask [2], building upon Mask R-CNN [1] , integrates YOLACT’s proton network and FOCS feature fusion mechanism to achieve finer instance segmentation.
+4 Experiment
+4.1 Experimental Settings
+Based on the requirements of brain tumor segmentation, this study conducts comparative experiments using two instance segmentation models: Mask R-CNN He et al.
+(2017) and BlendMask Chen et al.
+(2020).
+Mask R-CNN He et al.
+(2017) , as a classic two-stage method, employs a ResNet50-FPN backbone to generate region proposals and predict masks.
+BlendMask Chen et al.
+(2020), building upon Mask R-CNN He et al.
+(2017) , integrates YOLACT’s proton network and FOCS feature fusion mechanism to achieve finer instance segmentation.
 All models adopt a unified ResNet50-FPN backbone.
@@
 We apply random horizontal flipping as the basic data augmentation strategy.
-IV-B Dataset
-Our study utilizes a brain tumor dataset [46] comprising 3,064 T1-weighted MRI images, including meningioma (708 images), glioma (1,426 images), and pituitary tumor (930 images).
+4.2 Dataset
+Our study utilizes a brain tumor dataset Cheng et al.
+(2015) comprising 3,064 T1-weighted MRI images, including meningioma (708 images), glioma (1,426 images), and pituitary tumor (930 images).
 For the purpose of this study, we combine these three tumor types into a single unified category labeled as “tumor”.
@@
   • Binary tumor mask
-For experimental purposes, we randomly partition the dataset [46] into:
+For experimental purposes, we randomly partition the dataset Cheng et al. (2015) into:
   • Training set: 2,040 images (converted to VOC format)
@@
 Additionally, we conduct ratio experiments by systematically varying the calibration-to-test set ratios from 9:1 to 1:9 (with a single random partition for each ratio).
-IV-C Hyper parameter
-IV-C1 Mask R-CNN Configuration
-For our Mask R-CNN [1] implementation, we adopt several key hyperparameters optimized for tumor instance segmentation.
+4.3 Hyper parameter
+4.3.1 Mask R-CNN Configuration
+For our Mask R-CNN He et al.
+(2017) implementation, we adopt several key hyperparameters optimized for tumor instance segmentation.
 The model uses a single foreground class (NUM_CLASSES=1) and trains with a batch size of 8 (IMS_PER_BATCH=8) to balance GPU memory constraints and training efficiency.
@@
 The learning rate undergoes reductions by a factor of 0.1 (GAMMA=0.1) at epochs 16 and 22 (LR_STEPS=[16,22]), following standard practice for this architecture.
-IV-C2 BlendMask Configuration
-In the BlendMask [2] implementation, we configure the network with similar design considerations but different optimization parameters.
+4.3.2 BlendMask Configuration
+In the BlendMask Chen et al.
+(2020) implementation, we configure the network with similar design considerations but different optimization parameters.
 The model processes a single output class (NUM_CLASSES=1) for tumor segmentation using a batch size of 8 (IMS_PER_BATCH=8).
@@
 These parameters reflect empirical optimizations specific to brain tumor segmentation tasks.
-IV-C3 Conformal Prediction Framework
+4.3.3 Conformal Prediction Framework
 […]
-IV-D Evaluation Metrics
+4.4 Evaluation Metrics
 We employ the following quantitative measures to assess model performance:
-IV-D1 False Discovery Rate (FDR)
+4.4.1 False Discovery Rate (FDR)
 $\mathrm{FDR}=1-\frac{|\mathcal{P}\cap\mathcal{G}|}{|\mathcal{P}|}$ (11)
 where $\mathcal{P}$ denotes the predicted mask and $\mathcal{G}$ represents the ground truth. FDR quantifies the proportion of false positive predictions.
-IV-D2 False Negative Rate (FNR)
+4.4.2 False Negative Rate (FNR)
 $\mathrm{FNR}=1-\frac{|\mathcal{P}\cap\mathcal{G}|}{|\mathcal{G}|}$ (12)
 FNR measures the fraction of missed tumor regions in ground truth.
-IV-D3 Conformal Prediction Metrics
+4.4.3 Conformal Prediction Metrics
 For uncertainty quantification, we use:
@@
 Visual results present error curves comparing empirical FDR/FNR against theoretical bounds, while ablation studies show bar plots of error rates across different data splits (mean $\pm$ std over 10 trials).
-IV-E Theoretical Formulas
+4.5 Theoretical Formulas
 Theorem 1: FDR Control Formula
@@
 $\widehat{\lambda}=\inf\left\{\lambda:\frac{1}{n}\sum_{i=1}^{n}L_{i}(\lambda)\leq\alpha-\frac{B-\alpha}{n}\right\}.$ (16)
-IV-F Practical Steps
+4.6 Practical Steps
 Experiment 1 (Precision)
@@
 This choice ensures that the output of the prediction model remains as accurate as possible, avoiding excessive false positives (false discoveries), even if it might miss some target regions.
-We use both Mask R-CNN [1] and BlendMask [2] models and generate 10 test results through cross-validation.
+We use both Mask R-CNN He et al.
+(2017) and BlendMask Chen et al.
+(2020) models and generate 10 test results through cross-validation.
 As shown in Fig.
@@
 This choice ensures that the prediction model captures all possible target regions comprehensively, even at the cost of increasing some false positives (false negatives).
-To verify the effectiveness of this method, we use both Mask R-CNN [1] and BlendMask [2] models and generate 10 test results through cross-validation.
+To verify the effectiveness of this method, we use both Mask R-CNN He et al.
+(2017) and BlendMask Chen et al.
+(2020) models and generate 10 test results through cross-validation.
 As shown in Fig.
@@
 The ablation study results show that our method maintains strong robustness and achieves effective error rate control under different data distribution conditions.
-V Conclusion
+5 Conclusion
 In this work, we present a novel approach to brain tumor segmentation by integrating conformal prediction with advanced deep learning models.